OpenStudy (anonymous):

3/5p-4=26 solve and check.

OpenStudy (anonymous):

p=50

OpenStudy (anonymous):

can you please show me the work. Because I got 15 and i dont know why though.

OpenStudy (anonymous):

3/5p=30 ==> p=(30*5)/3 = 50

OpenStudy (anonymous):

(3/5)p = 30 ==> p=30*(5/3) =50

jhonyy9 (jhonyy9):

(3/5)p-4=26 3p/5=30 3p=150 p=150/3 p=50

jhonyy9 (jhonyy9):

so if you have thought no like how i have wrote it please us parantheses